OpenStudy (rosales0228):

NEED HELP PLEASE!!!! GIVING FAN AND MEDAL!!! Tina want to save money for school. Tina invest 1,000 in an account that pays and interest of 8%. How many years will it take for the account to reach 1,300 round your answer to the nearest hundredth

OpenStudy (rosales0228):

@cj49

OpenStudy (rosales0228):

@aaronq

OpenStudy (rosales0228):

@Sionainn

OpenStudy (anonymous):

Simple interest formula: I=PRT In this case, I=300 P=1000 R=.08 T=x What you need to do is substitute the values into the equation and solve for x.

OpenStudy (rosales0228):

thats just thing i suck at algebra and dont know how to do that

OpenStudy (anonymous):

Ok so the equation you have to solve is 300=1000*.08*x Simplify (1000*.08)... 300=80*x Simplfy (Divide by 80)... 3.75 = x It would take 3.75 years to get to $1300, assuming there are no deposits or withdrawals.

OpenStudy (rosales0228):

it was wrong:(

OpenStudy (anonymous):

Do you know if it's compound interest, not simple interest?

OpenStudy (rosales0228):

no just what i posted is all it says

jimthompson5910 (jim_thompson5910):

Assuming simple interest, the formula you'll use is A = P*(1+r*t) in this case A = 1300 P = 1000 r = 0.08 t = x if you were to use the formula I = P*r*t then you have to subtract 1300-1000 to get 300 to get the interest only. So I = 300, P = 1000, r = 0.08, t = x

jimthompson5910 (jim_thompson5910):

oh nvm, you have that already

OpenStudy (anonymous):

Yeah, I don't know why it wouldn't be correct

jimthompson5910 (jim_thompson5910):

the part "that pays and interest of 8%" looks like there's a typo there

jimthompson5910 (jim_thompson5910):

Rosales0228 can you post a screenshot of the whole problem?

OpenStudy (rosales0228):

it wont let me i copied and pasted the whole problem

jimthompson5910 (jim_thompson5910):

you're not able to use the printscreen function?

jimthompson5910 (jim_thompson5910):

try that out

OpenStudy (rosales0228):

A=p(1+r)^t p=amount of money invested r=interest rate % in decimal form

jimthompson5910 (jim_thompson5910):

so that implies they want to use compound interest

jimthompson5910 (jim_thompson5910):

A=p(1+r)^t 1300=1000(1+0.08)^x solve for x

OpenStudy (rosales0228):

@jim_thompson5910 ITS RIGHT!!!!! THANKS SOOOO MUCH!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome